If P1,000 accumulates to P1,500 when invested at a simple interest for three years, what is the annual rate of interest?
Answer by Theo(13342)

the formula for simple interest is i = p * r * n.

if p accumulates to 1500 from 1000, then i = 1500 - 1000 = 500.

when n = 3 years, the formula becomes:

500 = 1000 * r * 3

solve for r to get r = 500 / 1000 / 3 = .1667 per year rounded to 4 decimal places, or r = 16.67% per year rounded to 2 decimal places.

a more exact value would be i = 1/6 per year or r = (100/6)% per year.

take 1000 and multiply it by 1/6 * 3 and you get 1000 * 1/2 = 500.

that's your interest.

the annual rate of interest is (100/6)% per year, or 16.67% per year rounded to 2 decimal places.